If the price is $4 for a call option with strike of $30, what are the payoff and profit on a long position, if the option expires when the stock is $38.5?
Answer #1

Payoff = Max(0 , S - X)

= Max(0 , 38.5 - 30) = $8.5

Profit = Price of Underlying - Strike Price of Long Call - Premium Paid

= $38.5 - $30 - $4 = $4.5